OpenStudy (anonymous):

Solve the equation 1 over 5 x + 2 = 8 for x. 50 30 40 10

OpenStudy (anonymous):

\[\frac{1}{5}x+2=8\] \[\frac{1}{5}x=6\] \[x=6\times 5=30\]

OpenStudy (anonymous):